# A Comprehensive Survey on Continual Learning in Generative Models

### v1

@@ Abstract
The rapid advancement of generative models has enabled modern AI systems to comprehend and produce highly sophisticated content, even achieving human-level performance in specific domains. However, these models remain fundamentally constrained by catastrophic forgetting — a persistent challenge where adapting to new tasks typically leads to significant degradation in performance on previously learned tasks. To address this practical limitation, numerous approaches have been proposed to enhance the adaptability and scalability of generative models in real-world applications. In this work, we present a comprehensive survey of continual learning methods for mainstream generative models, including large language models, multimodal large language models, vision language action models, and diffusion models. Drawing inspiration from the memory mechanisms of the human brain, we systematically categorize these approaches into three paradigms: architecture-based, regularization-based, and replay-based methods, while elucidating their underlying methodologies and motivations. We further analyze continual learning setups for different generative models, including training objectives, benchmarks, and core backbones, offering deeper insights into the field. The project page of this paper is available at https://github.com/Ghy0501/Awesome-Continual-Learning-in-Generative-Models.
@@ 1 Introduction
The evolution of artificial intelligence is experiencing a paradigm shift from "understanding the world" to "creating the world". Traditional discriminative models, such as classification networks [1, 2, 3, 4, 5, 6] and object detectors [7, 8, 9, 10, 11], have achieved remarkable success over the past decade. These models are effective at learning decision boundaries from large-scale labeled datasets, enabling accurate recognition and judgment of known concepts. However, this discrimination-based intelligence has shown limitations in creativity and generalization. It excels at distinguishing known concepts within a single modality but struggles both to generate novel content and to establish semantic associations across multiple modalities. The rise of generative models addresses these limitations by employing large-scale pre-training to capture the underlying data distribution. As a result, mainstream generative models such as Large Language Models (LLMs) [12, 13, 14], Multimodal Large Language Models (MLLMs) [15, 16, 17, 18, 19], Vision-Language Action (VLA) [20, 21], and Diffusion Models [22, 23] are now capable of not only interpreting input features but also actively generating coherent text, realistic images, and even cross-modal content. As the field transitions from discriminative to generative paradigms, AI is evolving from merely recognizing the known to actively creating the new. This shift is redefining the foundations of human–machine collaboration by enabling more expressive, flexible, and multimodal generation capabilities.
Figure 1: An illustration of continual learning: human can acquire new knowledge throughout life without forgetting what was previously learned. Similarly, the goal of continual learning in generative AI is to endow models with the same ability—to continually learn without forgetting
Figure 2: A comparison of continual learning paradigm in discriminative models and generative models. Discriminative models typically operate within a single, fixed task format and modality. In contrast, generative models must address cross-modal Interactions, task-format inconsistencies, and more complex application demands.
@@
The success of generative models like GPT-4 fundamentally relies on large-scale pre-training and human alignment. However, this process is inherently dynamic, as language, data distributions, and user demands continually evolve. More critically, the impracticality of exhaustively covering all future scenarios necessitates models capable of continual learning (Figure 1), enabling adaptation to open-world environments [24, 25, 26]. For instance, fine-tuning general-domain models (e.g., ChatGPT) on specialized data (e.g., mathematics) often improves target performance at the cost of catastrophic forgetting [27, 28, 29], where prior knowledge is abruptly overwritten. When a model is trained sequentially on multiple downstream tasks, each new update further exacerbates catastrophic forgetting, leading to progressive deterioration in knowledge retention. A straightforward solution is to retain data from previous tasks and perform joint training when learning a new task. However, this approach is often impractical due to substantial storage requirements [30] or data privacy concerns [31]. In contrast, both children and adults rarely suffer from catastrophic forgetting when acquiring new knowledge [32, 33], which reveals a fundamental gap between current AI systems and human learning capabilities. To bridge this gap, the study of continual learning [34, 35, 36, 37, 38] has garnered increasing attention, aiming to equip models with the capability to acquire new knowledge sequentially while preserving previously learned tasks.
@@
Different from traditional discriminative models [39], continual learning in generative models presents greater challenges due to the complexity of generation objectives and the diversity of task forms. On the one hand, discriminative methods focus on modeling conditional probabilities (i.e. $P(y|x)$), where catastrophic forgetting primarily manifests as a drift in the decision boundary within the classification space [40, 41, 42, 43, 44]. In contrast, generative models aim to learn joint probability distributions (i.e. $P(x,y)$) or perform autoregressive generation, which requires continual learning to maintain both semantic coherence and knowledge completeness within an open-ended generative space. To achieve this, it is essential to refine the generative reasoning process alongside the underlying factual knowledge representation, thereby mitigating hallucinations [45, 46] and semantic inconsistencies [47]. On the other hand, generative models face highly diverse tasks during continual learning, such as visual question answering, visual captioning, and visual grounding in MLLMs. This variability increases the risk of catastrophic forgetting and places greater demands on multi-task knowledge integration. By comparison, discriminative models usually focus on a single task form, making continual learning comparatively simpler. Moreover, generative models involve multi-modal inputs and require effective cross-modal coordination, posing challenges that traditional continual learning methods, designed for unimodal tasks, cannot adequately address. These challenges, as illustrated in Figure 2, highlight that continual learning for generative models remains a promising yet underexplored research direction in the AI community.
@@
Existing continual learning surveys typically focus on specific generative models, such as Large Language Models (LLMs) [48, 49, 50, 51] or Multimodal Large Language Models (MLLMs) [52, 53, 54]. Although these works provide a systematic summary of continuous learning approaches within their respective domains, they often lack a broader discussion and connections across different generative models. Considering the increasing research interest in generative models and large-scale unified models within the AI community, we believe that an up-to-date and comprehensive survey will offer a holistic perspective to guide future work. In this paper, we present the first comprehensive review of continual learning for generative models, systematically analyzing key approaches and challenges across Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models. Furthermore, we unify these four types of generative models by exploring their shared methodological dimensions in the context of continual learning.
@@ 2 Overview of Continual Learning
(§ 7), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [VQACL [113], SGP [114], ProtoGroup [115], Lin et al. [116],
@@ 2.3 Taxonomy
Architecture-based methods mimic the brain’s modular organization [143, 144, 145, 146] through dynamic network expansion or modular design to isolate task-specific knowledge. These methods typically encapsulate such knowledge within distinct network modules. As model sizes grow, existing approaches increasingly adopt Parameter-Efficient Fine-Tuning (PEFT) techniques (e.g., LoRA [147] and Prompt tuning [148]) to store and adapt knowledge. Formally, let $\theta_{old}$ denote the model parameters that remain frozen during training on the current task. These typically include the backbone network and the extended sub-network associated with previously tasks. The parameters specific to the current task are denoted by $\theta_{new}$, which are usually the only trainable components. At this point, Eq (1) is rewritten as follows:
@@ 3.1.1 Training objectives
Figure 4: Representative architectures across different categories of continual learning in Large Language Models.
In continual learning scenario, the training data for the model arrives as a continuous stream $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{T}\}$, comprising a sequence of $T$ tasks. The objective of continual learning for LLM is to incrementally train on these $T$ datasets, aiming to maximize performance on the current task while preserving knowledge from previous tasks. Formally, at task $t$, the optimization objective in Eq. (13) can be reformulated as:
@@ 3.1.2 Backbone
  • BERT [163] employs a bidirectional encoder-only architecture based on the Transformer framework. Unlike autoregressive models, BERT leverages masked language modeling and next sentence prediction as its pretraining objectives, enabling it to learn contextualized word representations from both left and right contexts simultaneously. Each encoder layer comprises multi-head self-attention, feedforward networks, layer normalization, and residual connections.
  • T5 [164] adopts a standard encoder-decoder framework, where both the encoder and decoder are built upon the Transformer architecture. T5 reformulates all NLP tasks into a unified text-to-text format. Each layer in T5 consists of conventional multi-head attention and feedforward sublayers, with configurable embedding dimensions and attention heads depending on model scale.
  • GPT [165] series exemplifies a unidirectional decoder-only architecture, characterized by left-to-right autoregressive generation. GPT is composed of stacked decoder blocks, each containing masked multi-head attention, feedforward layers, and residual connections. The design is streamlined and restricts visibility to leftward context, making it well-suited for open-ended generation tasks. In most research settings, the GPT-2 model is used.
@@ 3.1.3 Benchmark
  • Long-sequence CL  [64] builds on the five datasets above by adding four tasks from the GLUE benchmark [167] (MNLI, QQP, RTE, SST-2), five tasks from SuperGLUE [168] (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB dataset [169], forming a longer continual learning evaluation sequence.
@@
  • TRACE [171] includes eight challenging datasets spanning domain-specific knowledge, multilingual ability, code generation, and mathematical reasoning. All datasets are uniformly formatted to allow automated evaluation. To measure forgetting in aligned LLMs during continual learning, TRACE introduces three core metrics: (1) General Ability Delta: tracks changes in performance on general tasks such as factual QA, reasoning, multilingual QA, commonsense reasoning, and reading comprehension; (2) Instruction Following Delta: measures shifts in instruction adherence on Self-instruct [172] and LIMA [173] prompt sets; (3) Safety Delta: evaluates changes in response safety using the CoNa dataset [174].
@@ 3.2.3 Prompt Expansion
As an alternative parameter-efficient fine-tuning approach for LLMs, prompt-based methods achieve continual learning by modularizing knowledge into prompts [190, 148], enabling efficient task adaptation and selective updates while reducing forgetting and interference. P-Prompts [64] learn a distinct prompt for each task and progressively concatenate them before feeding into a frozen backbone. It prevents forgetting by keeping task-specific prompts fixed after training, while enabling forward transfer through reuse of previous prompts. Q-Tuning [65] proposes a continual prompt tuning framework that maintains a prompt queue for old tasks and introduces new prompts for incoming tasks. It applies a low-rank adaptive knowledge aggregation mechanism for relevance reweighting, employs a PCA-based eviction rule to maintain constant queue size, and incorporates a globally shared prefix prompt with memory retention regularization to mitigate information loss across tasks. WAVE-CRE [66] proposes a continual relation extraction method that combines task-specific prompt pools with generative consolidation. The approach introduces a prompt pool for each task to capture intra-task variations and leverages a generative model to consolidate prior knowledge within shared parameters, effectively mitigating catastrophic forgetting without explicit data storage.
@@ 3.4.2 Generative Replay
Generative replay methods leverage LLMs to synthesize pseudo-samples from previous tasks, enabling memory consolidation without needing to store real historical data. For example, LAMOL [82] introduces a generation token to guide the model in producing pseudo-samples alongside new task learning. This approach enhances the model’s ability to retain old knowledge without incurring extra computational costs. HMI-LAMOL [83] proposes a brain-inspired generative replay mechanism, HMI, which uses BERT to encode training samples into compressed “memory traces.” These traces condition the generation of old task data, effectively alleviating sample imbalance and quality degradation, thus mitigating catastrophic forgetting in language models. Furthermore, PCLL [84] constructs natural language prompts for each task and employs a conditional variational autoencoder to model the task-specific latent distribution, generating high-quality pseudo-samples for replay. DCL [85] replaces the Gaussian distribution used in PCLL with a Dirichlet distribution to model the latent variables, enabling better representation of structural differences across tasks, and introduces Jensen–Shannon knowledge distillation at the logit layer to improve the model’s retention of old knowledge. SSR [86] further introduces a self-synthesizing reproduction framework, which utilizes the context learning ability of a base LLM to generate synthetic samples. These samples are refined by the latest fine-tuned model, and high-quality ones are selected via clustering algorithms, enabling the retention of prior knowledge without requiring historical data or additional generative models.
@@ 4 Continual Learning in Multimodal Large Language Models
Recent Multimodal Large Language Models [194], bulit upon advanced Large Language Models, have demonstrated impressive visual understanding capabilities by integrating visual encoders, cross-modal projector, and leveraging large-scale image-text corpus during training. Continual learning in MLLMs seeks to equip them with task-specific abilities across domains such as healthcare [195, 196], finance [197, 198], and autonomous driving [199, 200], without forgetting their previously learned knowledge. In contrast to discriminative models that predict output through classifiers, MLLMs generate responses by next-token-predict in an autoregressive way, making most continual learning methods designed for discriminative models less applicable. Additionally, compared to LLMs, MLLMs need to simultaneously process information from multiple modalities [201], thus posing a greater challenge.
In this section, we begin by introducing the setup for a MLLM continual learning task (Section 4.1), which includes training objectives, common model backbones and existing benchmarks. Subsequently, the existing literature is categorized into three strategies: architecture-based (Section 4.2), regularization-based (Section 4.3), and replay-based (Section 7) approaches.

### v2

@@ Abstract
The rapid advancement of generative models has enabled modern AI systems to comprehend and produce highly sophisticated content, even achieving human-level performance in specific domains. However, these models remain fundamentally constrained by catastrophic forgetting — a persistent challenge where adapting to new tasks typically leads to significant degradation in performance on previously learned tasks. To address this practical limitation, numerous approaches have been proposed to enhance the adaptability and scalability of generative models in real-world applications. In this work, we present a comprehensive survey of continual learning methods for mainstream generative models, encompassing large language models, multimodal large language models, vision-language action models, and diffusion models. Drawing inspiration from the memory mechanisms of the human brain, we systematically categorize these approaches into three paradigms: architecture-based, regularization-based, and replay-based methods, while elucidating their underlying methodologies and motivations. We further analyze continual learning setups for different generative models, including training objectives, benchmarks, and core backbones, offering deeper insights into the field. The project page of this paper is available at https://github.com/Ghy0501/Awesome-Continual-Learning-in-Generative-Models.
@@ 1 Introduction
The evolution of artificial intelligence is undergoing a paradigm shift from “understanding the world” to “creating the world”. Traditional discriminative models, such as classification networks [1, 2, 3, 4, 5, 6] and object detectors [7, 8, 9, 10, 11], have achieved remarkable success over the past decade. These models are effective at learning decision boundaries from large-scale labeled datasets, enabling accurate recognition and judgment of known concepts. However, this intelligence, based on discrimination, has shown limitations in creativity and generalization. It excels at distinguishing known concepts within a single modality but struggles both to generate novel content and to establish semantic associations across multiple modalities. The rise of generative models addresses these limitations by employing large-scale pre-training to capture the underlying data distribution. As a result, mainstream generative models such as Large Language Models (LLMs) [12, 13, 14], Multimodal Large Language Models (MLLMs) [15, 16, 17, 18, 19], Vision-Language Action (VLA) [20, 21], and Diffusion Models [22, 23] are now capable of not only interpreting input features but also actively generating coherent text, realistic images, and even cross-modal content. As the field transitions from discriminative to generative paradigms, AI is evolving from merely recognizing the known to actively creating the new. This shift is redefining the foundations of human-machine collaboration by enabling more flexible, and multimodal generation capabilities.
Figure 1: (a) Continual learning lets humans learn lifelong without forgetting; AI models aim to do the same while retaining old knowledge. (b) Discriminative models focus on perceptual tasks such as classification, detection, and segmentation, among others, whereas generative models additionally demand the capability to produce novel content grounded in perceptual understanding.
@@
The success of generative models like GPT-4o fundamentally relies on large-scale pre-training and human alignment. However, this process is inherently dynamic, as languages, data distributions, and user demands continually evolve. More importantly, it is infeasible to account for all possible future scenarios in advance, which underscores the need for models to possess continual learning capabilities comparable to those of humans, as illustrated in Figure 1 (a). Such capabilities are therefore essential for generative models to move beyond the closed-world assumption [24] and enable deployment in open-ended, real-world scenarios [25, 26, 27]. A central challenge in this context is catastrophic forgetting, where learning new information causes previously acquired knowledge to be overwritten [28, 29, 30]. For instance, fine-tuning general-domain models (e.g., ChatGPT) on specialized domains (e.g., mathematics) may improve in-domain performance at the expense of forgetting prior capabilities. When models are trained sequentially on multiple downstream tasks, each new update compounds this effect, leading to a progressive degradation in knowledge retention. A straightforward solution is to retain data from previous tasks and perform joint training with new data. However, this approach is often impractical due to significant storage costs [31] and data privacy concerns [32]. In contrast, both children and adults rarely experience catastrophic forgetting when acquiring new knowledge [33, 34], highlighting a fundamental gap between artificial and biological learning systems. To bridge this gap, the study of continual learning [35, 36, 37, 38, 39] has gained increasing attention, aiming to equip models with the ability to acquire knowledge sequentially while preserving previously learned skills.
@@
Compared to traditional discriminative models that focused on perceptual tasks, generative models in continual learning settings are expected to go a step further, as shown in Figure 1 (b), by generating new content based on understanding and recognition, owing to the diversity of modeling objectives and task formats. On the one hand, discriminative methods focus on modeling conditional probabilities (i.e. $P(y|x)$), where catastrophic forgetting primarily manifests as a drift in the decision boundary within the classification space [40, 41, 42, 43, 44]. In contrast, generative models aim to learn joint probability distributions (i.e. $P(x,y)$) or perform autoregressive generation, which requires continual learning to maintain both semantic coherence and knowledge completeness within an open-ended generative space. To achieve this, it is essential to refine the generative reasoning process alongside the underlying factual knowledge representation, thereby mitigating hallucinations [45, 46] and semantic inconsistencies [47]. On the other hand, generative models encounter highly diverse tasks in the continual learning process, with differences arising from factors such as new modality inputs, novel domain knowledge, and other evolving task characteristics. This variability increases the risk of catastrophic forgetting and places greater demands on multi-task knowledge integration. In contrast, discriminative models typically focus on a narrow range of tasks, such as classification, detection, or segmentation, which makes continual learning comparatively less challenging. Figure 2 demonstrates the continual learning scenarios across four representative generative models. Despite their impressive capabilities, enabling these models to learn continuously remains an open and pressing research challenge.
Figure 2: Illustrations of continual learning in generative models: Adapting to real-world demands by learning new instructions, knowledge, modalities, actions, and creations without catastrophic forgetting.
@@
Existing continual learning methods have been extensively developed for discriminative models [48, 49, 50, 51, 52]. In contrast, research on continual learning in generative models has generally focused on specific architectures, such as Large Language Models [53, 54, 55, 56] or Multimodal Large Language Models [57, 58, 59]. Although these works provide systematic overviews of continual learning techniques within their respective domains, they often lack broader discussions and cross-model connections. Motivated by this gap, and considering the growing interest in generative and large-scale unified models within the AI community, we believe that an up-to-date and comprehensive survey will offer a holistic perspective to guide future work. In this paper, we present the first comprehensive review of continual learning in mainstream generative models, systematically analyzing key approaches and challenges across Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models, as illustrated in Figure 2. In addition, we provide a unified perspective by identifying shared methodological dimensions of these generative models in the context of continual learning.
@@ 2 Overview of Continual Learning
(§ LABEL:MLLM:rep), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [VQACL [118], SGP [119], ProtoGroup [120], Lin et al. [121],
@@ 2.3 Taxonomy
Architecture-based methods mimic the brain’s modular organization [147, 148, 149, 150] through dynamic network expansion or modular design to isolate task-specific knowledge. These methods typically encapsulate such knowledge within distinct network modules. As model sizes grow, existing approaches increasingly adopt Parameter-Efficient Fine-Tuning (PEFT) techniques (e.g., LoRA [151] and Prompt tuning [152]) to store and adapt knowledge. Formally, let $\theta_{old}$ denote the model parameters that remain frozen during training on the current task. These typically include the backbone network and the extended sub-network associated with previous tasks. The parameters specific to the current task are denoted by $\theta_{new}$, which are usually the only trainable components. At this point, Eq (1) is rewritten as follows:
@@ 3.1.1 Training objectives
Figure 4: Overview of continual learning approaches in Large Language Models.
In a continual learning scenario, the training data for the model arrives as a continuous stream $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{T}\}$, comprising a sequence of $T$ tasks. The objective of continual learning for LLM is to incrementally train on these $T$ datasets, aiming to maximize performance on the current task while preserving knowledge from previous tasks. Formally, at task $t$, the optimization objective in Eq. (13) can be reformulated as:
@@ 3.1.2 Backbone
  • BERT [167] employs a bidirectional encoder-only architecture based on the Transformer framework. Unlike autoregressive models, BERT leverages masked language modeling and next-sentence prediction as its pretraining objectives, enabling it to learn contextualized word representations from both left and right contexts simultaneously. Each encoder layer comprises multi-head self-attention, feedforward networks, layer normalization, and residual connections.
  • T5 [168] adopts a standard encoder-decoder framework, where both the encoder and decoder are built upon the Transformer architecture. T5 reformulates all NLP tasks into a unified text-to-text format. Each layer in T5 consists of conventional multi-head attention and feedforward sublayers, with configurable embedding dimensions and attention heads depending on the model scale.
  • GPT [169] series exemplifies a unidirectional decoder-only architecture, characterized by left-to-right autoregressive generation. GPT is composed of stacked decoder blocks, each containing masked multi-head attention, feedforward layers, and residual connections. The design is streamlined and restricts visibility to a leftward context, making it well-suited for open-ended generation tasks. In most research settings, the GPT-2 model is used.
@@ 3.1.3 Benchmark
  • Long-sequence CL  [69] builds on the five datasets above by adding four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST-2), five tasks from SuperGLUE (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB dataset, forming a longer continual learning evaluation sequence.
@@
  • TRACE [172] includes eight challenging datasets spanning domain-specific knowledge, multilingual ability, code generation, and mathematical reasoning. All datasets are uniformly formatted to allow automated evaluation. To measure forgetting in aligned LLMs during continual learning, TRACE introduces three core metrics: (1) General Ability Delta: tracks changes in performance on general tasks such as factual QA, reasoning, multilingual QA, commonsense reasoning, and reading comprehension; (2) Instruction Following Delta: measures shifts in instruction adherence on Self-instruct and LIMA prompt sets; (3) Safety Delta: evaluates changes in response safety using the CoNa dataset.
@@ 3.2.3 Prompt Expansion
As an alternative parameter-efficient fine-tuning approach for LLMs, prompt-based methods achieve continual learning by modularizing knowledge into prompts [188, 152], enabling efficient task adaptation and selective updates while reducing forgetting and interference. P-Prompts [69] learn a distinct prompt for each task and progressively concatenate them before feeding into a frozen backbone. It prevents forgetting by keeping task-specific prompts fixed after training while enabling forward transfer through the reuse of previous prompts. Q-Tuning [70] proposes a continual prompt tuning framework that maintains a prompt queue for old tasks and introduces new prompts for incoming tasks. It applies a low-rank adaptive knowledge aggregation mechanism for relevance reweighting, employs a PCA-based eviction rule to maintain constant queue size, and incorporates a globally shared prefix prompt with memory retention regularization to mitigate information loss across tasks. WAVE-CRE [71] proposes a continual relation extraction method that combines task-specific prompt pools with generative consolidation. The approach introduces a prompt pool for each task to capture intra-task variations and leverages a generative model to consolidate prior knowledge within shared parameters, effectively mitigating catastrophic forgetting without explicit data storage.
@@ 3.4.2 Generative Replay
Generative replay methods leverage LLMs to synthesize pseudo-samples from previous tasks, enabling memory consolidation without needing to store real historical data. For example, LAMOL [87] introduces a generation token to guide the model in producing pseudo-samples alongside new task learning. This approach enhances the model’s ability to retain old knowledge without incurring extra computational costs. HMI-LAMOL [88] proposes a brain-inspired generative replay mechanism, HMI, which uses BERT to encode training samples into compressed “memory traces”. These traces condition the generation of old task data, effectively alleviating sample imbalance and quality degradation, thus mitigating catastrophic forgetting in language models. Furthermore, PCLL [89] constructs natural language prompts for each task and employs a conditional variational autoencoder to model the task-specific latent distribution, generating high-quality pseudo-samples for replay. DCL [90] replaces the Gaussian distribution used in PCLL with a Dirichlet distribution to model the latent variables, enabling better representation of structural differences across tasks, and introduces Jensen–Shannon knowledge distillation at the logit layer to improve the model’s retention of old knowledge. SSR [91] further introduces a self-synthesizing reproduction framework, which utilizes the context-learning ability of a base LLM to generate synthetic samples. These samples are refined by the latest fine-tuned model, and high-quality ones are selected via clustering algorithms, enabling the retention of prior knowledge without requiring historical data or additional generative models.
@@ 4 Continual Learning in Multimodal Large Language Models
Recent Multimodal Large Language Models [192], built upon advanced Large Language Models, have demonstrated impressive visual understanding capabilities by integrating visual encoders, cross-modal projectors, and leveraging large-scale image-text corpus during training. Continual learning in MLLMs seeks to equip them with task-specific abilities across domains such as healthcare [193, 194], finance [195, 196], and autonomous driving [197, 198], without forgetting their previously learned knowledge. In contrast to discriminative models that predict output through classifiers, MLLMs generate responses by next-token prediction in an autoregressive way, making most continual learning methods designed for discriminative models less applicable. Additionally, compared to LLMs, MLLMs need to simultaneously process information from multiple modalities [199], thus posing a greater challenge.
In this section, we begin by introducing the setup for an MLLM continual learning task (Section 4.1), which includes training objectives, common model backbones, and existing benchmarks. Subsequently, the existing literature is categorized into three strategies: architecture-based (Section 4.2), regularization-based (Section 4.3), and replay-based (Section 4.4) approaches.